OpenStudy (anonymous):

Which equation is in point-slope form for the given point and slope Point: (-2,4);slope:5

OpenStudy (gabebae):

y+y1= m (x-x1) (-2, 4) (-x, y) m= the slope y-4= 5(x+2)

OpenStudy (anonymous):

Thank you @gabebae